Your client is deciding between two investment choices: one that pays $100 per year in perpetuity,
and another that pays $100 per year for 100 years. The current market interest rate for investments
of similar risk is at 10% p.a. What is the present value of these two investments? Are they similar?
Explain.
1
Expert's answer
2020-09-08T15:42:45-0400

"PV= \sum^{\infty}_{k=1}" "\frac {C}{(1+r)^k}=\frac {C}{r}, >0"


Where PV= Present value

C=Cash flow

r= discount rate.


"PV= \frac {100}{0.1}=" $ 1000


PV of Annuity


"PV= \sum^{n}_{k=1}" "\frac {C}{(1+r)^k}" = "C" "\frac {1-(1+r)^ {-n}}{r}, r>0"


Where:

PV= Present value

C=Cash flow

r= discount rate.

n= number of periods


"PV= 100." "\frac {1-(1+0.1)^{-100}}{0.1}= 999.93<1000"


Thus: $999.93<$1000.


The present value of a perpetuity is greater than the present value of annuity.